1. To determine the number of deer in a game preserve, a conservationist catches 484 deer, tags them and lets them loose. Later, 618 deer are caught; 206 of them are tagged. How many deer are in the preserve?


2. Jack usually mows his lawn in 6 hours. Marilyn can mow the same yard in 7 hours. How much time would it take for them to mow the lawn together?

3. Benz bicycles 7 km/h faster than praedup. In the same time it takes praedup to bicycle 48 km, Benz can bicycle 69 km. how fast does each bicyclist travel?
Praedup travels what km/h?
Benz travels what km/h?

4. Find the LCM of 5(r-7) and 35(r-7)

Answer by stanbon(75887)

To determine the number of deer in a game preserve, a conservationist catches 484 deer, tags them and lets them loose. Later, 618 deer are caught; 206 of them are tagged. How many deer are in the preserve?
-----
Total population/caught : x/484
Total Tagged population/tagged caught: 618/206
----------
Proportion Equation:
x/484 = 618/206
x = 1452 (estimate of size of total population)
========================

2. Jack usually mows his lawn in 6 hours. Marilyn can mow the same yard in 7 hours. How much time would it take for them to mow the lawn together?
---------
Jack rate + Marilyn rate = together rate
1/6 + 1/7 = 1/x
(13/42) = 1/x
13x = 42
x = 42/13 = 3.23 hours to do the job together
==============
3. Benz bicycles 7 km/h faster than praedup.
In the same time it takes praedup to bicycle 48 km, Benz can bicycle 69 km. how fast does each bicyclist travel?
-----------------------
Benz DATA :
distance = 69 km ; rate = x+7 kph ; time = d/r = 69/(x+7)
----------------
Praedup DATA:
distance = 48 km ; rate = x kph ; time = d/r = 48/x
----------------
EQUATION:
time = time
69/(x+7) = 48/x
69x = 48x + 7*48
21x = 7*48
x = 16 km/h (Praedup speed)
x+7 = 23 km/h (Benz speed)
================================
4. Find the LCM of 5(r-7) and 35(r-7)
LCM = 35*(r-7)
===============
Cheers,
Stan H.